US. $               Currency Per U.S.
            Equivalent.                           $

UK. Pound (£) 1.5713              .6364
Canada dollar 1.0181.            .9822

a.Which would you rather have $100 or £100?
b. Which would you rather have $100 Canadian or £100?
c.What is the cross rate for Canadian dollars in terms of British pounds?)
d.What is cross-rate for British pounds in terms of Canadian dollars?

a.Rather have ____
b.Rather have ____
c. Cross rate C$_____/£
d. Cross rate £______/C$
Answer #1

Answer:

a. Rather have £100

Explanation:

Given:

UK. Pound (£) = $1.5713

Hence:

£100 = 100 * 1.5713 = $157.13

Hence you would rather have £100

b. Rather have C$100

Explanation:

Given:

Canada dollar = $1.0181

Hence:

C$100 = 100 * 1.0181 = $101.81

Hence you would rather have C$100

c. Cross rate C$1.5434 /£

Working:

Currency Per U.S. = £.6364

Currency Per U.S.= C$.9822

Cross rate per £ = 0.9822 /0.6364 = C$1.543369

d. Cross rate £0.6479 /C$

Working:

per £ = C$1.543369

Hence:

Per C$ = £ 1/1.543369= 0.6479